I live 260 miles from a popular mountain retreat. On my way there to do some mountain biking, my car had a break down forcing me to bike the rest of the way. If I drove 2 hours longer than I biked and average 60 mph driving and 10 mph biking, how many hours did I spend biking to the resort?
Answer by mananth(16946)

let biking distance be x
biking speed = 10 mph
driving distance = 260-x
driving speed = 60 mph
x/10 +2=(260-x)/60
x/10 -(260-x)/60 = -2
LCD = 60
multiply equation by 60
6x-260+x=-2*60
7x= -120+260
7x=140
/7
x= 20 miles
time biking = distance biking/speed biking
=20/10 = 2 hours